OpenStudy (jbo11):

I'm having trouble taking the indefinite integral of -1/(x^2-6x+9) dx Why isn't the answer -lnabs(x-3)+c

OpenStudy (johnweldon1993):

Hmm well lets see Firstly lets factor that bottom part \[-\int\limits^{}_{}\frac{1}{(x - 3)^2}dx \] Now make u = (x - 3) du = dx \[ -\int\limits^{}_{}\frac{1}{u^2}du \] The integral of 1/u^2 = -1/u + C so \[\ -\frac{-1}{u} + c \] \[\ \frac{1}{u} + C \] Plug back in u = (x - 3) \[\frac{1}{x - 3}+ C\]

OpenStudy (jbo11):

oooh ok thank you so much!

OpenStudy (johnweldon1993):

No problem!